a grade of 81 in test in which the mean is 75 and standard deviation is 12, corresponds to a z-score of
Answer #1

Here we have given that,

a data point(grade)=81

mean =E(x)= 75

and standard deviation(SD) = 12

we know that,

Z statistic is,

Z=(X-E(X))/SD

therefore,

Z=(81-75)/12

=6/12

Z=0.5